OpenStudy (anonymous):

find the solution of the system y=x^2+5 y=2x+4

OpenStudy (anonymous):

same thing as before yes?

OpenStudy (anonymous):

x^2+5 =2x+4 x^2-2x+1 (x-1)(x-1) ok? x=1

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

would it look like this http://go.hrw.com/math/midma/gradecontent/manipulatives/GraphCalc/graphCalc.html

OpenStudy (anonymous):

@fontez , pretty much did most of the work for you... all you need to do now is find the y-coordinate of the solution... plug in that x value into one of those equations to get the y-value.

OpenStudy (anonymous):

i cant read it

OpenStudy (anonymous):

ok i have the choices (1,6) and (-1,6) (-1,6) (1,6) (2,9)

OpenStudy (anonymous):

you'll need to learn this sometime... why not now? plug in x=1 into any one of those equations in your system to get the y-value.

OpenStudy (anonymous):

i dont understand how they got x=1

OpenStudy (dumbcow):

x^2+5 =2x+4 x^2-2x+1 =0 (x-1)(x-1) =0 x=1 which step do you get lost on?

OpenStudy (anonymous):

the beginning i dont understand why they are set equal to each other

OpenStudy (dumbcow):

because they both equal y....you are substituting (replacing) 2x+4 in for y

OpenStudy (anonymous):

oh okay then what next?

OpenStudy (dumbcow):

because its a quadratic, you need to set one side equal to 0

OpenStudy (anonymous):

ok so move one over to the other side to make one side 0?

OpenStudy (dumbcow):

right, what they did was subtract 2x and subtract 4 to other side

OpenStudy (anonymous):

x^2-2x+1=0 right?

OpenStudy (dumbcow):

yep, then next step they factored...its like backwards FOIL what factors of 1 add up to -2 ?

OpenStudy (dumbcow):

you could also use the quadratic formula here if you get stuck factoring

OpenStudy (anonymous):

how would i do the quadratic way?

OpenStudy (dumbcow):

\[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\] for ax^2 +bx +c = 0

OpenStudy (anonymous):

then i plug in the values correct?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

ok \[-(-2)\pm \sqrt{-2^2-4(1)(1)}\div2(1)\] like that?

OpenStudy (dumbcow):

yes, make sure you square the negative as well....b^2 will always be positive

OpenStudy (anonymous):

\[2\pm \sqrt{4-4} \div 2\]

OpenStudy (anonymous):

like that ?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

ok \[2\pm \sqrt{0}\div2\] how is that right

OpenStudy (dumbcow):

the 0 goes away --> 2/2 = 1

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

then i plug in the 1

OpenStudy (dumbcow):

yep, and it doesn't matter which equation

OpenStudy (anonymous):

y=6

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

thank you so much

OpenStudy (dumbcow):

your welcome